OpenStudy (anonymous):

Consider two straight lines L1 (y = 2x +1) and L2 (y=x). a) What is the slope of line L1 in respect to line L2? b) What is the equation of line L1 in respect to line L2? NOTE. Think that the original xy co-ordinate system has been rotated in respect to the origin so that line L2 defines the new x-axis x´ and y´ is the new y-axis!

OpenStudy (anonymous):

I need a smart person to answer this, please, I REALLY need to understand this

OpenStudy (amistre64):

a smart person eh..... hmmm

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

the slope of L1 with respect to L2 suggests to me that it is a relative answer. If L2 were flat, what would the slope of L1 be with it..

OpenStudy (amistre64):

y=x has a slope of 1, which is a tan(45)

OpenStudy (anonymous):

ok, I know this, but where do I go from here?

OpenStudy (amistre64):

y = 2x... has a slope of 2....

OpenStudy (anonymous):

so it is 2 :)

OpenStudy (amistre64):

i aint determined that yet :)

OpenStudy (amistre64):

it is the angle of slope 2 - slope1 but you use the "angles" and not the slopes

OpenStudy (amistre64):

what angle has a tan^-1(2)?

OpenStudy (anonymous):

oh, so, basically, the slope of L1 in relation to slope L2 will be 1, right?

OpenStudy (anonymous):

I am really bad at this sorry... forgot all the high school stuff

OpenStudy (amistre64):

tan(63.43) = 2 not really, since we have rotated the axis by a certain degree, we need to determine the degrees involved, then you trig stuff to re dress the problem

OpenStudy (amistre64):

.... we dont need the angles perse, just their ration....

OpenStudy (anonymous):

isn't there a simple way?

OpenStudy (amistre64):

cos(a-b) = cos(a)cos(b)-sin(a)sin(b)

OpenStudy (amistre64):

cos(a) = 1/sqrt(2) cos(b) = 1/sqrt(5) sin(a) = 1/sqrt(2) sin(a) = 2/sqrt(5)

OpenStudy (amistre64):

cos(a-b)=1/sqrt(2) * 1/sqrt(5) - 1/sqrt(2) * 2/sqrt(5)

OpenStudy (amistre64):

1/10 - 2/10 = -1/10

OpenStudy (amistre64):

might be better to work it in tans :) that way we get a new "slope"....tan = slope by the way

OpenStudy (amistre64):

tan(a-b) = tan(a)-tan(b) ----------- if I recall correctly 1+tan(a)tan(b)

OpenStudy (amistre64):

tan(a) = 1/1 tan(b)=2/1

OpenStudy (amistre64):

hold up, make a the bigger one... tan(a) = 2; tan(b)=1 2-1 ---- = 1/3 right? 1+2

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

the new slope is 1/3 :) what was the second part?

OpenStudy (anonymous):

equation of the line L1 in relation to eq of the line of L2

OpenStudy (amistre64):

y=(1/3)x + b just gotta determine where the lines intersect and use that as a parameter for the new system....

OpenStudy (amistre64):

they originally intersect at x=-1; y=-1

OpenStudy (amistre64):

the new point of intersection is the length of -sqrt(2)....

OpenStudy (amistre64):

so the new coord to fill in is (-sqrt(2),0)

OpenStudy (anonymous):

why sqrt(2)?

OpenStudy (amistre64):

0 = (1/3)(-sqrt(2)) + b b= sqrt(2)/3 y = (1/3)x - (sqrt(2)/3) should be it

OpenStudy (amistre64):

why sqrt(2)? the original lines cross at (-1,-1) when we rotate the axis to accomodate the new axises.... that point is still there, but we use the distance of it from the origin as the x intercept

OpenStudy (anonymous):

oh, ok, got it

OpenStudy (amistre64):

1-1-sqrt(2) is the triangle of a 45 degree..

OpenStudy (anonymous):

how did you get so good at maths?

OpenStudy (amistre64):

by messing up alot lol

OpenStudy (amistre64):

my son is nagging me for the laptop....ciao :)